To borrow $1,450, you are offered an add on interest loan at 8.5 percent with 12 monthly payments. Compute the 12 equal payments. (Round your answer to 2 decimal places.) 


Use the amount you borrowed and the monthly payments you computed to calculate the APR of the loan. Then, use that APR to compute the EAR of the loan. (Do not round intermediate calculations and round your answer to 2 decimal places.) 
Answer #1

Answer:

APR Calculations:

Interest APRT) = Principal ncipal * 100

123.2 APRT) = 1.450 * * 100

APR(r)=8.50%

EAR Calculations:

EAR(i) = (1+

0.085 EAR(i) = (1+2

EAR(i)=0.088391×100=8.84%

Interest APRT) = Principal ncipal * 100

123.2 APRT) = 1.450 * * 100

EAR(i) = (1+

0.085 EAR(i) = (1+2